OpenStudy (anonymous):

I'm having issues proving that for every n >= 4, n! >= 2^n using induction

zepdrix (zepdrix):

\[n=4 \rightarrow 4! \ge 2^4\]\[1*2*3*4 \ge 2*2*2*2\]\[24 \ge 16\] If n=k \[k! \ge 2^k\]\[1*2*3*4*....*(k-1)*(k) \ge 2_1 *2_2 *2_3* 2_4* ... *2_{k-1} *2_k\] Blahhh I don't remember induction. It seems like, from here, you can conclude that since the product of the first 4 terms is greater than 2^4, and EVERY other term is greater than 2, the remaining product will be greater than 2^k. I dunno sorry I forget how to set these up :P Hopefully someone can help ya..

zepdrix (zepdrix):

@TuringTest

OpenStudy (turingtest):

\[k!\ge2^k\implies\frac{2^k}{k!}\le1\]now for \(n=k+1\) we have\[(k+1)!\ge2^{k+1}\]\[k!(k+1)\ge2(2^k)\]\[k+1\ge2\left(\frac{2^k}{k!}\right)\]now because we know that \[\frac{2^k}{k!}\le1\]we know that\[2\left(\frac{2^k}{k!}\right)\le2\]and we know that k+1 is at least 5 since k is at least 4, so\[ k+1>2\ge2\left(\frac{2^k}{k!}\right)\]I think that is a solid proof